please help me write this into an equaton: four times the difference of a number and one is equal to six times the sum of the numbber and three.
Found 2 solutions by stanbon, ReadingBoosters:
Answer by stanbon(75887)

four times the difference of a number and one is equal to six times the sum of the numbber and three.
-----
Let the number be "x":
Equation:
4(x-1) = 6(x+3)
----
4x-4 = 6x+18
2x = -22
x = -11
-----------------
Cheers,
Stan H.

Answer by ReadingBoosters(3246)

4(x-1)
=
6(x+3)
OR
4(x-1) = 6(x+3)
..................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.